write equation in slope-intercept form
y-4=3/4(x+8)

Answer by lwsshak3(11628)

write equation in slope-intercept form
y-4=3/4(x+8)
y-4=(3/4)x+6
y=(3/4)x+10
This is an equation of a line with slope=3/4 and y-intercept=10